OpenStudy (cmtboy2):

Solve. 8 – 3x – x = 12 x =

OpenStudy (anonymous):

Well, firts you need to get all the x on one side and the numbers on the other one

OpenStudy (anonymous):

Calculate all the x's together and do the same with the numbers

OpenStudy (cmtboy2):

whats x

OpenStudy (anonymous):

didn't you write 8-3x-x=12 ?

OpenStudy (cmtboy2):

ya

OpenStudy (anonymous):

it's an unknown number, the one you need to find so that the calculation is true

OpenStudy (anonymous):

x =1

OpenStudy (anonymous):

4x = -4

OpenStudy (anonymous):

8 - -4 = 12

OpenStudy (anonymous):

bam

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

no, it's -1

OpenStudy (anonymous):

yea sorry

OpenStudy (cmtboy2):

im going with @L-Lawliet-L

OpenStudy (anonymous):

I just realized that

OpenStudy (anonymous):

since it is negative 4 for 4x

OpenStudy (cmtboy2):

7. (3 pt) Solve. 4x – 6 = −3x + 8 x = Question Resources

OpenStudy (anonymous):

yea I made a mistake with calculating it

OpenStudy (anonymous):

x =2

OpenStudy (anonymous):

here you do exactly the same

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

nope it doesn't = 2

OpenStudy (anonymous):

I do 3x on both sides

OpenStudy (anonymous):

add 3x on both sides and also add 6 on both sides and then calculate

OpenStudy (anonymous):

making it = 1.75

OpenStudy (anonymous):

right?

OpenStudy (cmtboy2):

Seth paints houses for a fee of $39 and an hourly rate of $13. Malcolm paints houses for a fee of $55 and an hourly rate of $11. In how many hours will Seth and Malcolm earn the same amount?

OpenStudy (anonymous):

oh wait I was right first time

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

I need to write this stuff down

OpenStudy (anonymous):

I am tired

OpenStudy (cmtboy2):

lol

OpenStudy (cmtboy2):

Seth paints houses for a fee of $39 and an hourly rate of $13. Malcolm paints houses for a fee of $55 and an hourly rate of $11. In how many hours will Seth and Malcolm earn the same amount?

OpenStudy (anonymous):

thank you for reposting

OpenStudy (anonymous):

39+13x=55+11x

OpenStudy (anonymous):

x is the time

OpenStudy (anonymous):

you know that seth is being paid 39 dollars and every hours he will get 13 more

OpenStudy (anonymous):

and malcolm is being paid 55 dollars and every hour he's getting 11 more

OpenStudy (anonymous):

what you need to find is the time

OpenStudy (cmtboy2):

13x11

OpenStudy (anonymous):

no, 13x-11x

OpenStudy (anonymous):

you take all the x's on the same side

OpenStudy (cmtboy2):

oh

OpenStudy (anonymous):

try the rest

OpenStudy (cmtboy2):

13x-11x=2x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and now take the numbers without an x on the other side and calculate them together

OpenStudy (cmtboy2):

?

OpenStudy (anonymous):

well now you have 39+2x =55

OpenStudy (cmtboy2):

55 is the answer

OpenStudy (anonymous):

no

OpenStudy (cmtboy2):

right

OpenStudy (anonymous):

take the 39 on the other side of the =

OpenStudy (cmtboy2):

im cunfused

OpenStudy (anonymous):

well, it's quite simple, now you have 39+2x=55, now you need to substract 39 on both sides

OpenStudy (cmtboy2):

39-2=37

OpenStudy (cmtboy2):

39-55=-16

OpenStudy (anonymous):

no, you got it wrong, you can't calculate a number with an x with a number without x

OpenStudy (cmtboy2):

oh sorry

OpenStudy (anonymous):

you substract 39 on both sides : 39-39+2x=55-39

OpenStudy (cmtboy2):

39-39=0 39-55=-16

OpenStudy (anonymous):

no, 55-39, not 39-55

OpenStudy (cmtboy2):

16

OpenStudy (anonymous):

and then divide both side by two so you dont have 2x anymore but just x left

OpenStudy (cmtboy2):

so divid 55by 2 then 39 by 2

OpenStudy (anonymous):

no, 16 by two

OpenStudy (cmtboy2):

16/2=8 and 2/16=0.125

OpenStudy (cmtboy2):

k now what

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

simply divide 2 by two and 16 by two

OpenStudy (cmtboy2):

0

OpenStudy (cmtboy2):

right

OpenStudy (anonymous):

the answer is 8

OpenStudy (mathmale):

Solve. 8 – 3x – x = 12 Combine the two "x" terms: Solve. 8 – 4x = 12 Subtract 8 from both sides: Solve. 8 – 4x = 12 -8 -8 -------------- -4x = 4 Divide both sides by -4. What is x? Check your answer by substituting it into the original equation.